OpenStudy (anonymous):

What is the length of the conjugate axis? (x-2)^2/(36) - (y+1)^2/(64) =1

OpenStudy (amistre64):

had to chk to be sure; the number in the denom of your - term; sqrt it, what do we get?

OpenStudy (amistre64):

if they are going to call them a transverse and a conjugate axis they might as well define the generic as: \[\frac{x^2}{t^2}-\frac{y^2}{c^2}=1\] where the c^2 is always under the - part

OpenStudy (anonymous):

so you're referring to C^2 which happens to be 64 in the equation

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2 * 8 = 16